OpenStudy (anonymous):

can someone help me with this plese? -3y=-6-21x x being -2

OpenStudy (anonymous):

just put the value of x in eq: -3y=-6-21*(-2) calculate....

OpenStudy (anonymous):

I did but I am not certain if I am doing it correctly -3y=-6-21*-2 -3y=-6+42 -3y=-6+6 0r -21 -6 stumped

OpenStudy (anonymous):

the answer should be... y= -12 i solved for y with the given x. you just plug it in for x and solve. its easy once you get it:)

OpenStudy (anonymous):

i should be.... -3y=-6-21(-2) than it becomes... -3y=-6+42 -3y=36 divide both sides by -3 and than you get y by itself and your answer